# Genesis-Metallicity: Universal Non-Parametric Gas-Phase Metallicity Estimation

### v2

@@ Abstract
1 Introduction
The “direct-method” provides a highly reliable measure of the gas-phase metallicity in galaxies. However, this method relies on an estimate of the electron temperature before the ionic abundances can be derived from the abundance-sensitive emission lines. Unfortunately, the often-faint temperature-sensitive emission lines such as [O III] and [O II] remain mostly elusive in large spectroscopic surveys, hindering the application of the direct-method on large samples. This has made the “strong” emission lines such as the [O II] and [O III] doublets the most commonly used proxies for the gas-phase metallicities of galaxies with available rest-optical spectroscopy (see, e.g., Savaglio et al., 2005; Erb et al., 2006; Maiolino et al., 2008; Mannucci et al., 2009; Zahid et al., 2011, 2014; Wuyts et al., 2012, 2016; Belli et al., 2013; Henry et al., 2013; Kulas et al., 2013; Cullen et al., 2014; Yabe et al., 2014; Maier et al., 2014; Steidel et al., 2014; Troncoso et al., 2014; Kacprzak et al., 2016, 2015; Sanders et al., 2015, 2021; Hunt et al., 2016; Onodera et al., 2016; Suzuki et al., 2017; Curti et al., 2017, 2024a; Langeroodi et al., 2023; Langeroodi & Hjorth, 2023; Heintz et al., 2023; Nakajima et al., 2023; Chemerynska et al., 2024; Sarkar et al., 2024). This practice is commonly known as the “strong-line” metallicity estimation: several polynomial relations between various strong-line ratios and gas-phase metallicity are calibrated either empirically on samples with direct-method measurements (see, e.g., McGaugh, 1991; Pilyugin et al., 2010; Pilyugin & Grebel, 2016; Curti et al., 2017; Jiang et al., 2019; Nakajima et al., 2022) or against the predictions of photoionization models (see, e.g., McCall et al., 1985; Denicoló et al., 2002; Kewley & Dopita, 2002; Hirschmann et al., 2023).
@@ 1 Introduction
Third; recent parametric calibrations at high redshifts based on NIRSpec spectroscopy indicate noticeable deviations from the local-universe calibrations (Sanders et al., 2024; Laseter et al., 2024), potentially suggesting a non-universality in the strong-line method. However, as shown by Nakajima et al. (2018) and Nakajima et al. (2022), the 2D-projected relationships between the line ratios and gas-phase metallicity are influenced by the ionization parameter. Therefore, high-redshift deviations from the locally-calibrated parametric strong-line estimators are expected, as the high-redshift galaxies exhibit systematically higher ionization parameters. This is evidenced by their observed extremely high O32, EW(H), and EW(H) values (Langeroodi et al., 2023; Langeroodi & Hjorth, 2024; Rinaldi et al., 2023), indicative of high ionization parameters (Kewley & Dopita, 2002; Hirschmann et al., 2023) and bursty star formation histories (Smit et al., 2016; Langeroodi & Hjorth, 2024). Nonetheless, it is essential for any strong-line calibration to capture such dependencies and remain insensitive to these systematics.
@@
Here, we overcome these caveats by developing a “non-parametric” strong-line metallicity estimator. We achieve this by a kernel density estimate (KDE; Silverman, 1986; Scott, 1992) of the probability density function (PDF) in the multi-dimensional space of emission line observables and gas-phase metallicity (Section 4). This PDF is then used to estimate the gas-phase metallicity for any combination of input emission line observables. We calibrate our strong-line estimator on a sample of 1510 galaxies at with direct-method metallicity measurements, the largest of such compilations to date (Sections 2 and 3). In particular, we report 122 new direct-method metallicity measurements at based on NIRSpec multi-shutter assembly (MSA; Jakobsen et al., 2022; Ferruit et al., 2022) spectroscopy; this corresponds to a fold increase in the sample size of directly-measured metallicities. We show that the O2, O3, and EW(H) measurements are sufficient for a gas-phase metallicity estimate that is more accurate than 0.09 dex. Our calibration is universal, meaning that its accuracy does not depend on the target redshift. We make genesis-metallicity (Langeroodi, 2025) and its calibration data available at https://github.com/langeroodi/genesis_metallicity (the calibration data will be made available upon publication).
2 Data
@@
In this Section, we present an overview of the spectra utilized in our strong-line metallicity calibration. This data consists of 1510 spectra with direct-method metallicity measurements, including 171 galaxies observed with the NIRSpec MSA, 122 of which are reported for the first time in this work and the rest are taken from the literature (Section 2.1); 1213 galaxies observed with ground-based instruments (Section 2.2); and 126 high-metallicity spectra generated by stacking the SDSS spectra (Section 2.3). Figure 1 provides an overview of this sample. The line fluxes are reported in Table 1. We note that the H, H, H, and H Balmer lines are used for dust reddening correction of emission lines. For this purpose, we assumed a Calzetti et al. (2000) dust curve111Several recent studies have found evidence that while non-negligible dust is already present in early-universe galaxies, their attenuation curves might deviate from those in local universe (Langeroodi et al., 2024; Ciesla et al., 2024; Burgarella et al., 2024, 2025; Markov et al., 2025a, b; McKinney et al., 2025; Fisher et al., 2025). As such, a dedicated investigation of potential systematic biases caused by adopting the Calzetti et al. (2000) or other similar dust curves in this and other similar work is timely. However, this goes beyond the scope of our current study. and case-B recombination222The dust attenuation module is available at https://github.com/langeroodi/genesis_metallicity.
2.1 NIRSpec
@@
We searched the JADES DR3 (D’Eugenio et al., 2024) NIRSpec MSA medium-resolution333Due to its low spectral resolution, the prism grating almost never resolves the [O III] line from H. Exceptions can occur at , where these lines fall at relatively high-resolution () prism wavelengths (Williams et al., 2023; Schaerer et al., 2024; Curti et al., 2024b). spectra for [O III] and [O II] detections. For this purpose, we used pPXF (Cappellari & Emsellem, 2004; Cappellari, 2017, 2022) to measure the emission line fluxes. For the objects covered in multiple JADES observations, we stacked the spectra from repeated gratings to enhance the signal. We adopted the spectroscopic redshifts reported by the JADES team as a starting point, and for each object ran pPXF on the medium-resolution spectra covering its [O III] and [O II] emission. We then visually inspected the subsample with either [O III] or [O II] flux signal-to-noise ratios (S/N) greater than 3. We confirm 138 galaxies with robust [O III] detections (), 20 of which also exhibit robust [O II] detection (). These exclude the confirmed broad-line AGN from Maiolino et al. (2024). We also fitted the prism spectra of the [O III]-detected sample to achieve full coverage of the [O II]; H; [O III]; H; [O III]; H; and [O II] lines. The NIRSpec medium-resolution spectra as well as the corresponding best-fit pPXF models for some example galaxies with [O III] and [O II] detections are presented in Appendix A and Figures 7, 8, 9, 10, 11, 12, 13, and 14.
@@ 2.2 Ground-based
2.3 SDSS stacks
Andrews & Martini (2013) and Curti et al. (2017) showed that the individual SDSS spectra can be stacked to enhance the [O III] signal and enable direct-method metallicity measurements for the less-explored high-metallicity () region of the parameter space. Employing a similar approach, we selected 58207 non-AGN spectra from the MPA-JHU catalog with [O II]; H; [O III]; and H detections.
We stacked these spectra on a three-dimensional grid of reddening-corrected O2, O3, and EW(H). This is in contrast with Curti et al. (2017), where the spectra are stacked on a 2-dimensional grid of reddening-corrected O2 and O3. We chose the 3-dimensional grid because our strong-method calibration relies on O2, O3, and EW(H) to estimate the gas-phase metallicity (see Section 4). We binned the O2 axis in 0.1 dex intervals, the O3 axis in 0.1 dex intervals, and the EW(H) axis in 1 dex intervals. We stacked the spectra using the stacking algorithm detailed in Langeroodi & Hjorth (2024). We used pPXF to measure the line fluxes and EW(H) for the stacks. As reported by Curti et al. (2017), the [Fe II] line is a common source of systematic offsets in [O III] flux measurement of very high metallicity galaxies. To avoid such systematics, we add the [Fe II] line to the list of emission lines fitted by pPXF. We identified 126 stacks with robust [O III] detections (), 108 of which also exhibit significant [O II] detections ().
3 Direct Measurements
We measure the ionic oxygen abundances and gas-phase metallicities (O/H) by modelling the emission lines with a 2-zone H II region (Stasińska, 1982; Garnett, 1992). This corresponds to a bithermal nebula model, where the low-ionization zone containing species such as and the high-ionization zone containing species such as are traced by different temperatures. Assuming an electron density (), temperature-sensitive line ratios can be used to calculate the electron temperature of each zone. In turn, these temperature measurements allow to derive the ionic abundances of each zone from the abundance-sensitive line fluxes. Where available, we use the [O II] and [S II] lines to estimate the electron densities (adopting the Kisielius et al. 2009 and Tayal & Zatsarinny 2010 collision strengths, respectively), while assuming otherwise. The derived temperatures and abundances are only weakly sensitive to the assumed electron density at the density regimes common for galaxies (see, e.g., Curti et al., 2017; Nakajima et al., 2023; Isobe et al., 2023). We describe the electron temperature measurements in Section 3.1 and ionic abundances and gas-phase metallicity measurements in Section 3.2. These measurements are reported in Table 1.
3.1 Electron temperatures
@@
3.2 Metallicities
We calculate the ionic abundances from the [O III]/H line ratios, employing the getIonAbundance routine of PyNeb and assuming the (O III) electron temperatures calculated in Section 3.1. Similarly, the ionic abundances are calculated from the [O II]/H line ratios, assuming the (O II) electron temperatures calculated in Section 3.1. Whenever there is no (O II) measurement available, we used the (O II) estimator calibrated in Section 3.1 to estimate the (O II) based on the measured O2, O3, EW(H), and (O III). We assume that the and are the most abundant oxygen ions, and derive the oxygen abundances (gas-phase metallicity) as the sum of and ionic abundances.
@@ 3.2 Metallicities
4 Strong-line calibration
We use the distribution of the calibration data in the 4-dimensional space of O2, O3, EW(H), and gas-phase metallicity for a non-parametric calibration of a strong-line metallicity estimator. Figure 4 shows 4 classic projections of the data, frequently used for the parametric calibration of the strong-line metallicity estimators. We adapt a method similar to that described in Section 3.1 for the non-parametric calibration. In brief, we use a kernel density estimate (KDE) in the 4-dimensional space of O2, O3, EW(H), and gas-phase metallicity to estimate the probability density function (PDF) non-parametrically based on the distribution of the calibration data (our multivariate probability density estimation is described in more detail in Appendix C.1; real-valued 5D images of this probability density estimate are available online in machine-readable format). This PDF is then used to estimate the gas-phase metallicity for any combination of input O2, O3, and EW(H). The marginalization procedure is described in detail in Section 3.1. In this calibration, we only include the subsample with direct-method metallicity uncertainties lower than 0.2 dex.
We evaluate the accuracy of our strong-line metallicity estimator with a leave-one-out cross-validation approach, similar to that described in Section 3.1 (additional validity tests are provided in Appendix C.2). In each iteration, we exclude one data point from the calibration sample, calibrate the metallicity estimator on the remaining data, and use this estimator to estimate the metallicity of the excluded point based on its O2, O3, and EW(H) measurements. Figure 5 shows the strong-line gas-phase metallicity estimates vs. those measured by the direct method. Our metallicity estimations are more accurate than 0.09 dex, defined as the absolute strong-line vs. direct metallicity offset which contains of the estimates.
@@ 4 Strong-line calibration
As shown in Figure 6, EW(H) and O32 are tightly correlated. Since O32 is widely accepted as an ionization parameter estimator (see, e.g., Kewley & Dopita, 2002; Hirschmann et al., 2023), this correlation suggests that EW(H) closely traces the ionization parameter as well. Photoionization models imply that the relationship between the strong-line ratios and gas-phase metallicity is sensitive to the ionization parameter (Nakajima et al., 2018, 2022). As such, including an ionization parameter estimator such as O32 or EW(H) in the strong-line metallicity calibration is expected to increase its accuracy. Indeed, Nakajima et al. (2022) showed that the parametric strong-line metallicity calibration can be improved, particularly at the low-metallicity end, by splitting calibration into three ionization branches as traced by the observed EW(H).
The tight correlation between the EW(H) and O32 (see Figure 6) might suggest that O2 and O3 should be sufficient to calibrate an optimal strong-line metallicity estimator; i.e., suggesting that including the EW(H) information is unnecessary. This is because EW(H) is accurately predicted from its tight correlation with O32, and the O32 information is already captured by including the O2 and O3. We test this by removing the EW(H) axis from our strong-line metallicity calibration, and using a KDE in the 3-dimensional space of O2, O3, and gas-phase metallicity. This slightly yet noticeably decreases the accuracy of our metallicity estimator to 0.17 dex. Hence, EW(H) is providing additional information beyond what is captured by O2 and O3. This seems intuitive from Figure 6, where the offset from the average EW(H)-O3 relation seems to correlate with both the gas-phase metallicity and redshift, as indicated by the color-coding in the left and right panels, respectively. Similarly, we find that including the EW(H) information slightly improves the accuracy of our (O II) estimator (Section 3.1).

### v3

@@ Abstract
I Introduction
The “direct-method” provides a highly reliable measure of the gas-phase metallicity in galaxies. However, this method relies on an estimate of the electron temperature before the ionic abundances can be derived from the abundance-sensitive emission lines. Unfortunately, the often-faint temperature-sensitive emission lines such as [O III] and [O II] remain mostly elusive in large spectroscopic surveys, hindering the application of the direct-method on large samples. This has made the “strong” emission lines such as the [O II] and [O III] doublets the most commonly used proxies for the gas-phase metallicities of galaxies with available rest-optical spectroscopy (see, e.g., Savaglio et al., 2005; Erb et al., 2006; Maiolino et al., 2008; Mannucci et al., 2009; Zahid et al., 2011, 2014; Wuyts et al., 2012, 2016; Belli et al., 2013; Henry et al., 2013; Kulas et al., 2013; Cullen et al., 2014; Yabe et al., 2014; Maier et al., 2014; Steidel et al., 2014; Troncoso et al., 2014; Kacprzak et al., 2016, 2015; Sanders et al., 2015, 2021; Hunt et al., 2016; Onodera et al., 2016; Suzuki et al., 2017; Curti et al., 2017, 2024a; Langeroodi et al., 2023; Langeroodi and Hjorth, 2023; Heintz et al., 2023; Nakajima et al., 2023; Chemerynska et al., 2024; Sarkar et al., 2024). This practice is commonly known as the “strong-line” metallicity estimation: several polynomial relations between various strong-line ratios and gas-phase metallicity are calibrated either empirically on samples with direct-method measurements (see, e.g., McGaugh, 1991; Pilyugin et al., 2010; Pilyugin and Grebel, 2016; Curti et al., 2017; Jiang et al., 2019; Nakajima et al., 2022) or against the predictions of photoionization models (see, e.g., McCall et al., 1985; Denicoló et al., 2002; Kewley and Dopita, 2002; Hirschmann et al., 2023).
@@ I Introduction
Third; recent parametric calibrations at high redshifts based on NIRSpec spectroscopy indicate noticeable deviations from the local-universe calibrations (Sanders et al., 2024; Laseter et al., 2024), potentially suggesting a non-universality in the strong-line method. However, as shown by Nakajima et al. (2018) and Nakajima et al. (2022), the 2D-projected relationships between the line ratios and gas-phase metallicity are influenced by the ionization parameter. Therefore, high-redshift deviations from the locally-calibrated parametric strong-line estimators are expected, as the high-redshift galaxies exhibit systematically higher ionization parameters. This is evidenced by their observed extremely high O32, EW(H), and EW(H) values (Langeroodi et al., 2023; Langeroodi and Hjorth, 2024; Rinaldi et al., 2023), indicative of high ionization parameters (Kewley and Dopita, 2002; Hirschmann et al., 2023) and bursty star formation histories (Smit et al., 2016; Langeroodi and Hjorth, 2024). Nonetheless, it is essential for any strong-line calibration to capture such dependencies and remain insensitive to these systematics.
@@
Here, we overcome these caveats by developing a “non-parametric” strong-line metallicity estimator. We achieve this by a kernel density estimate (KDE; Silverman, 1986; Scott, 1992) of the probability density function (PDF) in the multi-dimensional space of emission line observables and gas-phase metallicity (Section IV). This PDF is then used to estimate the gas-phase metallicity for any combination of input emission line observables. We calibrate our strong-line estimator on a sample of 1510 galaxies at with direct-method metallicity measurements, the largest of such compilations to date (Sections II and III). In particular, we report 122 new direct-method metallicity measurements at based on NIRSpec multi-shutter assembly (MSA; Jakobsen et al., 2022; Ferruit et al., 2022) spectroscopy; this corresponds to a fold increase in the sample size of directly-measured metallicities. We show that the O2, O3, and EW(H) measurements are sufficient for a gas-phase metallicity estimate that is more accurate than 0.09 dex. Our calibration is universal, meaning that its accuracy does not depend on the target redshift. We make genesis-metallicity (Langeroodi, 2025) and its calibration data available at https://github.com/langeroodi/genesis_metallicity.
II Data
@@
In this Section, we present an overview of the spectra utilized in our strong-line metallicity calibration. This data consists of 1510 spectra with direct-method metallicity measurements, including 171 galaxies observed with the NIRSpec MSA, 122 of which are reported for the first time in this work and the rest are taken from the literature (Section II.1); 1213 galaxies observed with ground-based instruments (Section II.2); and 126 high-metallicity spectra generated by stacking the SDSS spectra (Section II.3). Figure 1 provides an overview of this sample. The line fluxes are reported in Table 1. We note that the H, H, H, and H Balmer lines are used for dust reddening correction of emission lines. For this purpose, we assumed a Calzetti et al. (2000) dust curve111Several recent studies have found evidence that while non-negligible dust is already present in early-universe galaxies, their attenuation curves might deviate from those in local universe (Langeroodi et al., 2024; Ciesla et al., 2024; Burgarella et al., 2024, 2025; Markov et al., 2025b, a; McKinney et al., 2025; Fisher et al., 2025). As such, a dedicated investigation of potential systematic biases caused by adopting the Calzetti et al. (2000) or other similar dust curves in this and other similar work is timely. However, this goes beyond the scope of our current study. and case-B recombination222The dust attenuation module is available at https://github.com/langeroodi/genesis_metallicity.
II.1 NIRSpec
@@
We searched the JADES DR3 (D’Eugenio et al., 2024) NIRSpec MSA medium-resolution333Due to its low spectral resolution, the prism grating almost never resolves the [O III] line from H. Exceptions can occur at , where these lines fall at relatively high-resolution () prism wavelengths (Williams et al., 2023; Schaerer et al., 2024; Curti et al., 2024b). spectra for [O III] and [O II] detections. For this purpose, we used pPXF (Cappellari and Emsellem, 2004; Cappellari, 2017, 2022) to measure the emission line fluxes. For the objects covered in multiple JADES observations, we stacked the spectra from repeated gratings to enhance the signal. We adopted the spectroscopic redshifts reported by the JADES team as a starting point, and for each object ran pPXF on the medium-resolution spectra covering its [O III] and [O II] emission. We then visually inspected the subsample with either [O III] or [O II] flux signal-to-noise ratios (S/N) greater than 3. We confirm 138 galaxies with robust [O III] detections (), 20 of which also exhibit robust [O II] detection (). These exclude the confirmed broad-line AGN from Maiolino et al. (2024). We also fitted the prism spectra of the [O III]-detected sample to achieve full coverage of the [O II]; H; [O III]; H; [O III]; H; and [O II] lines. The NIRSpec medium-resolution spectra as well as the corresponding best-fit pPXF models for some example galaxies with [O III] and [O II] detections are presented in Appendix A and Figures 7, 8, 9, 10, 11, 12, 13, and 14.
@@ II.2 Ground-based
II.3 SDSS stacks
Andrews and Martini (2013) and Curti et al. (2017) showed that the individual SDSS spectra can be stacked to enhance the [O III] signal and enable direct-method metallicity measurements for the less-explored high-metallicity () region of the parameter space. Employing a similar approach, we selected 58207 non-AGN spectra from the MPA-JHU catalog with [O II]; H; [O III]; and H detections.
We stacked these spectra on a three-dimensional grid of reddening-corrected O2, O3, and EW(H). This is in contrast with Curti et al. (2017), where the spectra are stacked on a 2-dimensional grid of reddening-corrected O2 and O3. We chose the 3-dimensional grid because our strong-method calibration relies on O2, O3, and EW(H) to estimate the gas-phase metallicity (see Section IV). We binned the O2 axis in 0.1 dex intervals, the O3 axis in 0.1 dex intervals, and the EW(H) axis in 1 dex intervals. We stacked the spectra using the stacking algorithm detailed in Langeroodi and Hjorth (2024). We used pPXF to measure the line fluxes and EW(H) for the stacks. As reported by Curti et al. (2017), the [Fe II] line is a common source of systematic offsets in [O III] flux measurement of very high metallicity galaxies. To avoid such systematics, we add the [Fe II] line to the list of emission lines fitted by pPXF. We identified 126 stacks with robust [O III] detections (), 108 of which also exhibit significant [O II] detections ().
III Direct Measurements
We measure the ionic oxygen abundances and gas-phase metallicities (O/H) by modelling the emission lines with a 2-zone H II region (Stasińska, 1982; Garnett, 1992). This corresponds to a bithermal nebula model, where the low-ionization zone containing species such as and the high-ionization zone containing species such as are traced by different temperatures. Assuming an electron density (), temperature-sensitive line ratios can be used to calculate the electron temperature of each zone. In turn, these temperature measurements allow to derive the ionic abundances of each zone from the abundance-sensitive line fluxes. Where available, we use the [O II] and [S II] lines to estimate the electron densities (adopting the Kisielius et al. 2009 and Tayal and Zatsarinny 2010 collision strengths, respectively), while assuming otherwise. The derived temperatures and abundances are only weakly sensitive to the assumed electron density at the density regimes common for galaxies (see, e.g., Curti et al., 2017; Nakajima et al., 2023; Isobe et al., 2023). We describe the electron temperature measurements in Section III.1 and ionic abundances and gas-phase metallicity measurements in Section III.2. These measurements are reported in Table 1.
III.1 Electron temperatures
@@
III.2 Metallicities
We calculate the ionic abundances from the [O III]/H line ratios, employing the getIonAbundance routine of PyNeb and assuming the (O III) electron temperatures calculated in Section III.1. Similarly, the ionic abundances are calculated from the [O II]/H line ratios, assuming the (O II) electron temperatures calculated in Section III.1. Whenever there is no (O II) measurement available, we used the (O II) estimator calibrated in Section III.1 to estimate the (O II) based on the measured O2, O3, EW(H), and (O III). We assume that the and are the most abundant oxygen ions, and derive the oxygen abundances (gas-phase metallicity) as the sum of and ionic abundances.
@@ III.2 Metallicities
IV Strong-line calibration
We use the distribution of the calibration data in the 4-dimensional space of O2, O3, EW(H), and gas-phase metallicity for a non-parametric calibration of a strong-line metallicity estimator. Figure 4 shows 4 classic projections of the data, frequently used for the parametric calibration of the strong-line metallicity estimators. We adapt a method similar to that described in Section III.1 for the non-parametric calibration. In brief, we use a kernel density estimate (KDE) in the 4-dimensional space of O2, O3, EW(H), and gas-phase metallicity to estimate the probability density function (PDF) non-parametrically based on the distribution of the calibration data (our multivariate probability density estimation is described in more detail in Appendix C.1; real-valued 4D images of this probability density estimate are available online in machine-readable format). This PDF is then used to estimate the gas-phase metallicity for any combination of input O2, O3, and EW(H). The marginalization procedure is described in detail in Section III.1. In this calibration, we only include the subsample with direct-method metallicity uncertainties lower than 0.2 dex.
We evaluate the accuracy of our strong-line metallicity estimator with a leave-one-out cross-validation approach, similar to that described in Section III.1 (additional validity tests are provided in Appendix C.2). In each iteration, we exclude one data point from the calibration sample, calibrate the metallicity estimator on the remaining data, and use this estimator to estimate the metallicity of the excluded point based on its O2, O3, and EW(H) measurements. Figure 5 shows the strong-line gas-phase metallicity estimates vs. those measured by the direct method. Our metallicity estimations are more accurate than 0.09 dex, defined as the absolute strong-line vs. direct metallicity offset which contains of the estimates.
@@ IV Strong-line calibration
As shown in Figure 6, EW(H) and O32 are tightly correlated. Since O32 is widely accepted as an ionization parameter estimator (see, e.g., Kewley and Dopita, 2002; Hirschmann et al., 2023), this correlation suggests that EW(H) closely traces the ionization parameter as well. Photoionization models imply that the relationship between the strong-line ratios and gas-phase metallicity is sensitive to the ionization parameter (Nakajima et al., 2018, 2022). As such, including an ionization parameter estimator such as O32 or EW(H) in the strong-line metallicity calibration is expected to increase its accuracy. Indeed, Nakajima et al. (2022) showed that the parametric strong-line metallicity calibration can be improved, particularly at the low-metallicity end, by splitting calibration into three ionization branches as traced by the observed EW(H).
The tight correlation between the EW(H) and O32 (see Figure 6) might suggest that O2 and O3 should be sufficient to calibrate an optimal strong-line metallicity estimator; i.e., suggesting that including the EW(H) information is unnecessary. This is because EW(H) is accurately predicted from its tight correlation with O32, and the O32 information is already captured by including the O2 and O3. We test this by removing the EW(H) axis from our strong-line metallicity calibration, and using a KDE in the 3-dimensional space of O2, O3, and gas-phase metallicity. This slightly yet noticeably decreases the accuracy of our metallicity estimator to 0.17 dex. Hence, EW(H) is providing additional information beyond what is captured by O2 and O3. This seems intuitive from Figure 6, where the offset from the average EW(H)-O3 relation seems to correlate with both the gas-phase metallicity and redshift, as indicated by the color-coding in the left and right panels, respectively. Similarly, we find that including the EW(H) information slightly improves the accuracy of our (O II) estimator (Section III.1).